OpenStudy (anonymous):

Medal for help ty :) Which equation represents the line that passes through the points (6,-3)and (-4,-9)?

OpenStudy (anonymous):

Use the equation for slope: (y-k)/(x-h), with the points (x,y) and (h,k).

OpenStudy (haseeb96):

m=y2-y1/x2-x1 use this formuala here and first find slope that is m=3/5

OpenStudy (haseeb96):

now use the point slope formula here and put the slope and any one point in this formula y-y1 =m ( x -x1)

OpenStudy (anonymous):

y+3/x-6=(-9+3/-4-6 y+3/x-6=-6/-10 y+3/x-6=3/5 5(y+3)=3(x-6) =>5y+15=3x-18 =>5y=3x-18-15 =>5y=3x-33 =>y=3/5x-33/5

OpenStudy (anonymous):

its D

OpenStudy (haseeb96):

so here point is (6,-3)

OpenStudy (haseeb96):

D is the correct answer

OpenStudy (anonymous):

D my answer ?

OpenStudy (haseeb96):

y-6 = 3/5 (x +3 ) is the correct answer

OpenStudy (haseeb96):

GOOD JOB einestine

OpenStudy (anonymous):

Ty trying hard I have a deadline to hand in my work this coming week to pass ugg I really am grateful for your help :)

OpenStudy (haseeb96):

but in actually the @zaibali.qasmi is the honour of Medal so give him medal and remove the best response from my side